OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=sec(x)-2cos(x) at the point (pi/3,1).

OpenStudy (loser66):

Where are you stuck?

OpenStudy (anonymous):

i got the derv part which is sec(x)tan(x)+2sin(x)

OpenStudy (anonymous):

how do i put in the points

OpenStudy (loser66):

just plug pi/3 in

OpenStudy (loser66):

remember y' is slope of the line, and at pi/3, y' = something, not y. so that, ignore 1 from the point at this time.

OpenStudy (anonymous):

is this right y=sec(pi/3)tan(pi/3)+2sin(pi/3)

OpenStudy (loser66):

y' , not y , Again :)

OpenStudy (loser66):

and yes, that is. y' =??

OpenStudy (anonymous):

that where i dont get to solve it

OpenStudy (loser66):

\[y' = sec (\pi/3)tan(\pi/3)+2sin(\pi/3)\] sec (pi/3) = 1/cos (pi/3) =1/1/2 =2 tan (pi/3) = sqrt 3 sin pi/3 = sqrt3/2 just plug number in

OpenStudy (loser66):

I messed up tan (pi/3) , correct it

OpenStudy (loser66):

\[y'= 2(this ~is~ sec (pi/3) * \dfrac{\sqrt3}{3}(this~ is~ tan(pi/3) +2\dfrac{1}{2}(this~ is~ sin (pi/3))\]

OpenStudy (loser66):

y'= \(2\sqrt 3/3+1\) got it??

OpenStudy (anonymous):

oh okay thx

OpenStudy (anonymous):

its not right